Required information

[The following information applies to the questions displayed below.]

At the beginning of Year 1, Oak Consulting had the following normal balances in its accounts:

Account Balance
Cash $ 29,200
Accounts receivable 16,400
Accounts payable 12,200
Common stock 20,200
Retained earnings 13,200


The following events apply to Oak Consulting for Year 1:

  1. Provided $74,500 of services on account.
  2. Incurred $3,000 of operating expenses on account.
  3. Collected $50,000 of accounts receivable.
  4. Paid $36,000 cash for salaries expense.
  5. Paid $13,680 cash as a partial payment on accounts payable.
  6. Paid a $8,600 cash dividend to the stockholders.

c. Show the beginning balances and the events in a horizontal statements model such as the following one: (In the Statement of Cash Flows column, use the initials "OA" for operating activities, "FA" for financing activities and "NC" for net change in cash. Select "NA" wherever required. Enter any decreases to account balance and cash outflows with a minus sign.)

OAK CONSULTING
Effect of Transactions on the Financial Statements for Year 1
Balance Sheet Income Statement Statement of Cash Flows
Assets = Liabilities + Stockholders' Equity Revenue Expenses = Net Income
No Cash + Accounts Receivable = Accounts Payable + Common Stock + Retained Earnings
Bal. + = + + =
1 + = + + =
2 + = + + =
3 + = + + =
4 + = + + =
5 + = + + =
6 + = + + =
Bal. 0 + 0 = 0 + 0 + 0 0 0 = 0 0
Answer #1

c. Show the beginning balances and the events in a horizontal statements model such as the following one: (In the Statement of Cash Flows column, use the initials "OA" for operating activities, "FA" for financing activities and "NC" for net change in cash. Select "NA" wherever required. Enter any decreases to account balance and cash outflows with a minus sign.)

OAK CONSULTING
Effect of Transactions on the Financial Statements for Year 1
Balance Sheet Income Statement Statement of Cash Flows
Assets = Liabilities + Stockholders' Equity Revenue Expenses = Net Income
No Cash + Accounts Receivable = Accounts Payable + Common Stock + Retained Earnings
Bal. 29200 + 16400 = 12200 + 20200 + 13200 =
1 NA + 74500 = NA + NA + 74500 74500 NA = 74500 NA
2 NA + NA = 3000 + NA + -3000 NA 3000 = -3000 NA
3 50000 + -50000 = NA + NA + NA NA NA = NA 50000 OA
4 -36000 + NA = NA + NA + -36000 NA 36000 = -36000 -36000 OA
5 -13680 + NA = -13680 + NA + NA NA NA = NA -13680 OA
6 -8600 + NA = NA + NA + -8600 NA NA = NA -8600 FA
Bal. 20920 + 40900 = 1520 + 20200 + 40100 74500 39000 = 35500 -8280 NC